OpenStudy (anonymous):

help . K is the midpoint of . If = 3x - 4 and , find the length of

OpenStudy (anonymous):

. K is the midpoint of AB . If AK = 3x - 4 and , find the length of AB

OpenStudy (anonymous):

@AriPotta

OpenStudy (anonymous):

*and KB=1/2x+6

OpenStudy (anonymous):

K is the midpoint of AB . If AK = 3x - 4 and KB=1/2x+6 , find the length of AB

OpenStudy (anonymous):

@AriPotta

OpenStudy (anonymous):

@tkhunny

OpenStudy (aripotta):

i don't get it?

OpenStudy (anonymous):

this is the whole question @AriPotta K is the midpoint of AB . If AK = 3x - 4 and KB=1/2x+6 , find the length of AB

OpenStudy (aripotta):

AK and KB should have the slope

OpenStudy (aripotta):

shouldn't they?

OpenStudy (anonymous):

its lines and coordinates assignment

OpenStudy (aripotta):

it makes no sense though

OpenStudy (anonymous):

oh

OpenStudy (tkhunny):

A-----K----B m(AK) = 3x-4 m(KB) = (1/2)x + 6 Since K is the midpoint, 3x-4 = (1/2)x + 6 Is it trickier than that?

OpenStudy (aripotta):

oooh

OpenStudy (anonymous):

nvm i found an answer on google :)

OpenStudy (anonymous):

3x-4=1/2 x + 6 2.5x=10 x=4 AK=8 so AB =16